OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x+7 6x-2y = 12 A.There is no solution. B.(14, 12) C.(12, 14) D.There are an infinite number of solutions.

OpenStudy (anonymous):

Note that \[ 6x - 2y = 6x - 2(3x-7) = 6x - 6x + 14 = 14, \]which cannot be 12, so no solution.

OpenStudy (igreen):

Substitute 3x - 7 for y in the second equation: 6x - 2y = 12 6x - 2(3x - 7) = 12 Distribute -2 into the parenthesis: 6x - 6x + 14 = 12 Subtract 6x - 6x, what do you get? @ihateschool99

OpenStudy (igreen):

@Ahaanomegas Don't give out answers.

OpenStudy (anonymous):

Sorry; at least I showed my work.